i am trying to prepare to take the act . but i do not know how to sole this quistion:
(4z+3)(z-2)

Answer by stanbon(75887)

i am trying to prepare to take the act . but i do not know how to sole this quistion:
(4z+3)(z-2) = 0
----
Since the product is zero one or those factors must be zero.
----
Solve 4z+3 = 0
4z = -3
z = -3/4
------------------
Solve z-2 = 0
z = 2
-------------------
Final Answer: z = -3/4 or z = 2
=======================================
Cheers,
Stan H.